### v1

# Generative LiDAR Editing with Controllable Novel Object Layouts

Abstract
We propose a framework to edit real-world Lidar scans with novel object layouts while preserving a realistic background environment. Compared to the synthetic data generation frameworks where Lidar point clouds are generated from scratch, our framework focuses on new scenario generation in a given background environment, and our method also provides labels for the generated data. This approach ensures the generated data remains relevant to the specific environment, aiding both the development and the evaluation of algorithms in real-world scenarios. Compared with novel view synthesis, our framework allows the creation of counterfactual scenarios with significant changes in the object layout and does not rely on multi-frame optimization. In our framework, the object removal and insertion are supported by generative background inpainting and object point cloud completion, and the entire pipeline is built upon spherical voxelization, which realizes the correct Lidar projective geometry by construction. Experiments show that our framework generates realistic Lidar scans with object layout changes and benefits the development of Lidar-based self-driving systems.
I Introduction
Data is a key factor in the realization of self-driving. Collecting a huge amount of data with good diversity is necessary for understanding the corner cases, which are low-frequency and high-risk real-world situations. Both the development and the testing of self-driving systems could benefit from data with good long-tail coverages. However, the low-frequency nature means that it is inherently expensive to collect corner-case data, and as progress is being made on relatively easier and more frequent cases, capturing even rarer events only becomes an even more challenging task. That’s why the research community has a growing interest in data generation, which allows for acquiring new data at a lower cost and potentially higher diversity compared with real-world data collection.
Early efforts of data synthesis rely on gaming engines and graphical rendering. This approach generates data from highly configurable virtual environments, but the scalability is limited, and the domain gap to the real data is non-negligible. Recent progress in generative modeling makes end-to-end data generation possible, largely improving the scalability and realism. However, users generally have less control over the specifics of the generated data. Another notable trend is novel-view synthesis from real-world sequential observations, allowing the resimulation of a real-world scenario from different viewpoints, which helps analyze and improve the algorithms. However, the variation flexibility of the re-simulated data is limited by the visible signal.
In this paper, we propose to solve the task of scene editing with controllable novel object layouts, with a specific focus on Lidar scans—a widely used and reliable sensory modality in autonomous vehicles. This new task enables us to freely remove and insert objects with arbitrary poses in a real Lidar scan, not limited by the visible information. On the other hand, the generated data is associated with a real environment, building a more direct connection to real-world performance. An illustration of this task is shown in Fig. 1. Leveraging recent progress in conditional lidar data generation and object point cloud completion, we build a lightweight framework without scene reconstruction, neural rendering, or end-to-end generation from scratch. We bake the prior knowledge of Lidar projective geometry into the framework through spherical voxelization so that the occlusion relationship and the scanning pattern are correct by construction. Experiments show that our framework enables flexible, controllable, and realistic Lidar scene editing, which is beneficial for both the development and testing of self-driving algorithms.
II Related Work
We will review the literature related to Lidar data generation. Most work did not directly address the same problem as targeted in this paper, but a lot of them are relevant and prepared the techniques that are useful in our problem.
II-A Lidar simulation
Lidar simulation refers to the approach of building a 3D environment with 3D object assets and casting rays on the 3D environment to create point clouds through ray-surface intersection [1]. While graphics-based simulation engines [2, 3, 4] facilitate the rendering, the requirement of pre-built 3D assets limits the scalability, and the rendering and the quality of the 3D assets both contribute to the domain gap compared with real data. [5] constructed the 3D assets from real-world scans to avoid manual asset creation and used learning-based postprocessing to reduce the domain gap. Later work [6] further investigated various Lidar phenomena that could improve the fidelity of Lidar simulators.
II-B Lidar novel-view synthesis
The novel-view synthesis (NVS) of Lidar point clouds and Lidar simulation share the same conceptual procedure of 3D construction and rendering, but today’s Lidar NVS realizes these two steps in a unified learnable framework called neural rendering. The learned implicit geometry-appearance model and the differentiable rendering are both optimized towards reproducing more realistic synthesis, allowing a smaller domain gap and a more automatic workflow than traditional simulation [7]. [8] modeled Lidar effects like secondary returns and ray dropping in the neural rending framework. [9, 10] also leveraged image sequences for additional information. [11, 12] explicitly the temporal dimension to account for the moving objects. [13] rendered both Lidar scans and camera images from the neural field. These approaches realized the playback of a real-world scenario with changed viewpoints and object states, but the pose of the observer and the objects cannot deviate too much from reality, or the rendering quality will deteriorate.
II-C End-to-end Lidar data generation
End-to-end approaches bypass the two-step procedure of 3D construction and rendering. Inspired by the success of text- and image-based generative models, researchers proposed to generate Lidar scans through a fully-differentiable neural network. This approach produces data with promising diversity and realism but has limited controllability. [14] applied a diffusion model on equirectangular range-view images, leveraging the progress in image diffusion models. [15] also used range images but compressed the data into a latent space before diffusion. It also supported conditioning like semantic maps and RGB images. [16] applied a generative transformer framework MaskGIT [17] in BEV-voxelized point cloud generation and allowed basic editing by swapping the discrete latent codes. Recent work started to generate temporally consistent sequences of Lidar scans, incorporating the idea of world modeling [18]. [19] took traffic maps as input conditions, improving the controllability of Lidar sequence generation. The end-to-end generated point clouds generally do not come with corresponding labels, limiting their value in practical development.
II-D Lidar scan modification
Generative methods can also be used to apply modifications on an existing point cloud. For example, [20] used a transformer to upsample a sparse point cloud. [21] applied scene completion on a single Lidar scan to obtain a dense and complete scene. [22] allows inserting the same object in an RGB image and in a Lidar point cloud in a consistent manner. Overall, the editing flexibility is limited.
The task in this paper, Lidar scan editing with novel object layout, is related to but different from all the tasks discussed above. Our output is connected to a real scene, similar to the NVS task, but we do not rely on reconstruction, which allows more flexibility in the object manipulation. The task is generative but does not build scenes from scratch, emphasizing the controllability of object layouts.
III Problem Formulation
We address the problem of generating realistic synthetic Lidar data for autonomous driving. Rather than generating data completely from scratch, our approach builds upon real-world scenes, and we aim to manipulate the dynamic objects while preserving the background environment.
The static background of a Lidar scene, denoted as , consists of stationary elements such as streets, trees, buildings, and other fixed background. excludes any foreground dynamic objects of interest, for example, vehicles. The foreground objects in the scene are represented by . A typical driving scene, denoted as , includes both background and foreground objects: . Since it is not feasible to sense the entire from a single viewpoint, we define an observation function , which transforms the scene into a partial-view Lidar point cloud: , which represents the real sensor output for an autonomous vehicle.
The task of Lidar generative editing can be formally defined as follows. We assume that we have a dataset of real-world Lidar scans . Given a true Lidar scan , the goal is to generate a synthetic point cloud , where is the true foreground objects in the scan, and is a modified set of foreground objects. Here, may consist of the same objects as but with different poses, or it may contain entirely different objects. We assume that all possible objects in are observed in some .
We tackle this problem by decomposing it into three sub-problems: estimating , , and , corresponding to object removal, object point cloud completion, and object insertion. They are introduced in Sec. IV.
IV Method
IV-A Overview
Our scene editing framework is illustrated in Fig. 2. We first introduce the spherical voxel representation (Sec. IV-B), which is the foundation of our entire pipeline for the efficient modeling of the Lidar projective geometry. Then, we present the three components in our framework. The first involves removing unwanted objects from the Lidar scene (Sec. IV-C). Given an object to be removed, we identify the voxels occupied by the object’s point cloud. We mask these voxels and all voxels occluded by them with an inpainting mask. We remove points in the masked voxels and use a generative point cloud inpainting model to fill the background.
In the second stage, we collect an object library composed of dense full-shape point clouds for every object in all Lidar scans (Sec. IV-D), so that we have a wide range of objects to be inserted back into the Lidar scans with flexible poses. We use the segmentation masks within each Lidar scene to extract partial-view point clouds of vehicles. These partial views are then transformed into full point clouds using a learning-based point cloud completion method.
In the third stage, we insert new objects into the Lidar point cloud (Sec. IV-E). We start by selecting an object from the full-view object library, and specifying a desired pose for its placement. The dense point cloud of this object is then positioned on the ground at the specified pose. The resampling of the inserted objects and the occlusion are then both handled in the spherical voxel representation.
It is worth noting that in our framework, moving an object to a different pose is treated as both removing and re-adding the same object. This approach helps resolve any gaps in the background and allows for significant pose changes.
IV-B Spherical Voxelization
The spherical voxel representation is core to our entire pipeline for efficient occlusion handling and density control, as illustrated in Fig. 3. Spherical voxelization discretizes the space around the origin based on three parameters : radius , azimuth , and elevation . This approach closely mirrors how a LiDAR scan operates, where rays are emitted from the sensor (origin) at specific angles, allowing point resampling with a pattern consistent with real LiDAR sensors. Voxels sharing the same azimuth and elevation values resemble a single LiDAR ray, embedding the occlusion relationship into the data structure (see Fig. 3-right). While the range-view representation could also model the occlusion, the 2D projection loses the 3D structure. Preserving the 3D structure allows effective feature learning and convenient point cloud manipulation in our pipeline.
IV-C Object removal
The object removal has two steps: identifying the area affected by the removal of an object and generating background points in this area that are consistent with the surrounding environment. We call these two steps de-occlusion masking and background inpainting, respectively.
IV-C1 De-occlusion masking
Given the spherical voxelization of the input, a de-occlusion mask refers to a binary mask over the voxels, where the voxels occupied or occluded by the object to be removed are labeled positive. The occupancy of the masked voxels needs to be predicted because of the scene change (object removed) and the visibility change (occlusion removed). We first mask the voxels occupied by the object point cloud. Then, all voxels sharing the same azimuth and elevation indices with larger radius indices than the occupied voxels are also masked (see Fig. 3-right).
IV-C2 Background inpainting
The background inpainting can be viewed as a conditional generation problem. We recast the UltraLidar framework [16] for this task. UltraLidar can be viewed as a Lidar version of MaskGIT [17], which is a generative model for images with transformer-based multi-step autoregression in a discrete latent space. The training has two stages, as depicted in Fig. 4. First, we train a VQ-VAE [23] to encode raw point clouds into a latent feature map with discrete latent codes. We use a bird’s-eye-view (BEV) 2D latent map to represent the encoded point cloud. Notice that our BEV uses the azimuth-radius coordinate from the spherical voxelization, meaning that the voxels along the elevation axis are compressed to a single point in the latent feature map. It is slightly different from the physical top-down projection and also different from the Cartesian voxelization used in UltraLidar. We use the original point cloud scans to learn this encoder.
The second step is to learn a generative model in the latent space for background inpainting. Similar to MaskGIT [17], a bi-directional transformer is trained to map an incomplete BEV latent feature map to a complete latent feature map, which can be decoded to a point cloud. The feature prediction is cast as a classification problem because the feature space is discrete. The prediction is iterative. We use a mask to keep track of which tokens in the BEV feature map need to be predicted. In every iteration, the transformer predicts the masked tokens from the known ones, but only a subset of the predicted tokens with the top classification confidence is preserved, and they will be treated as known tokens in the next iteration. The other tokens remain masked and need to be predicted in the following iterations until the feature map is complete. The BEV mask is initialized from the voxel mask that is used to remove points from the input point cloud. In training, we apply the voxel mask on random object-free voxels, so that the transformer is trained to predict only the background for the masked tokens. At inference, we use the de-occlusion voxel mask introduced in Sec. IV-C1.
While we follow the UltraLidar framework [16], the object removal in UltraLidar is realized by replacing the token at the object position with a background discrete feature, which needs to be manually specified and does not explicitly inpaint the areas occluded by the removed objects. Our object-removal module resolves both problems.
IV-D Object library construction
We create a library of objects with full 3D shapes to be inserted at arbitrary poses in a Lidar scene. The objects are collected from all Lidar scans in our training data. We choose the point cloud as the shape representation of objects for two reasons. First, we can resample points from the shape conveniently in the spherical voxel representation and do not need expensive ray-casting on meshes or volumetric rendering on implicit fields. Second, point cloud completion techniques are well-established to facilitate full-shape generation. We applied a pretrained model of the state-of-the-art point cloud completion network, AnchorFormer [24], in our framework.
IV-E Object insertion
The insertion process has two major steps. First, given the desired pose to insert an object, we determine the height to make sure the object is placed on the ground. Second, we resolve the occlusion caused by object insertion, and resample the point cloud on the spherical voxel grids to yield a pattern consistent with the Lidar scan.
IV-E1 Object placement
We assume that a user only specifies the x-y coordinate and the yaw angle of the desired insertion pose, and the insertion algorithm needs to figure out the rest of the degrees of freedom based on the existing background point cloud. We further assume zero pitch and roll angles as are commonly assumed in self-driving datasets [25, 26]. Therefore, we are left with height to be determined.
To achieve this, we use the segmentation mask to find the ground points. Given a desired object position , we locate the nearest ground point in the x-y plane and align its value with that of the lowest point in the inserted object point cloud, allowing us to fully define the vehicle’s pose in the 3D space. Finally, we place the completed full point cloud of the target object into the Lidar scan.
IV-E2 Resampling and resolving occlusion
The point cloud, in its current state, presents two critical problems. First, the dense inserted point cloud does not match the surrounding environment in terms of scan pattern and density. Second, the occlusion relationships are incorrect.
These two problems can be efficiently fixed with our spherical voxelization technique. We locate the voxels occupied by the point cloud of the inserted vehicle. Then, we pick the center of each occupied voxel as the resampled points for the shape. Now, the inserted object has a pattern and density that are consistent with the overall Lidar scan. In the end, we resolve the occlusion. For each coordinate with at least one occupied voxel, an occlusion mask is applied to all voxels with radius , where is the smallest radius of the occupied voxels in the 1D ray. All the masked voxels are set to empty.
V Experiments and results
This paper aims at a novel Lidar editing task that is different from the literature. Therefore, it is not straightforward to apply a well-established evaluation protocol and compare with existing work. We will separate the experiments into two major components, i.e., object removal and object insertion, so that we can validate the effectiveness of each step and show that the overall framework provides a practical solution for Lidar scene editing.
We conduct experiments on the nuScenes-LidarSeg [25] dataset, which is composed of 40,000 point clouds sampled from 1,000 driving scenes. The point clouds are collected by a 32-beam Velodyne HDL32E Lidar. We use the official training and validation splits. All results are trained on the Lidar scans in the training set and evaluated using scans in the validation set. We focus on the manipulation of vehicle-type objects, including cars, trucks, and buses. The specific use of the dataset in each experiment is explained in the regarding sections.
V-A Object removal and background inpainting
The goal of this experiment is to test that our approach can inpaint the background points when objects are removed. The tricky thing is that we do not have the ground truth background behind an object. Therefore, we create artificial occlusion masks over the background area in a Lidar scan and ask our model to recover the masked area. In this way, we have the ground truth background to facilitate evaluation.
In order to make the artificial masks similar to the ones caused by the occlusion of real objects, we create the masks from the bounding boxes of objects in nuScenes. This process is illustrated in Fig. 5. We calculate the average size of the 3D bounding boxes of all vehicles in nuScenes. Then, we put the bounding box at a fixed distance, with the heading orthogonal to the viewing direction, so that the box occludes the most azimuth range. We rotate the bounding box around the ego vehicle till the bounding box falls into an object-free sector of the Lidar scan to be masked. We pick the distance as 10 meters to create a substantial occlusion.
We evaluate the performance of our background inpainting using statistical metrics and perceptual metrics. For the statistical metrics, we leverage the Maximum-Mean Discrepancy (MMD) and the Jensen-Shannon Divergence (JSD). Similar to [16], we construct a 2D histogram along the ground plane and use the voxel occupancy instead of the number of points for the bin count. However, our evaluation has two differences from that in [16]. First, our 2D histogram is in the azimuth-radius coordinate. Second, we manually set the count for all bins that are not in the generated area to zero and re-normalize the histogram to sum to one, because we only generate the point cloud for a masked area instead of generating the whole point cloud from scratch. For the perceptual metrics, we use the Frechet Sparse Volume Distance (FSVD) and the Frechet Point-based Volume Distance (FPVD). Following [15], FSVD is evaluated using MinkowskiNet [27], and FPVD is evaluated using SPVCNN [28], both pretrained on nuScenes.
| Perceptual | Statistical | |||
| Metrics | FSVD | FPVD | JSD | MMD |
| Baseline | 0.189 | 0.191 | 0.615 | 6.85 |
| Ours | 0.169 | 0.173 | 0.584 | 6.61 |
| Metrics | mAP | Car | Truck | Bus |
| Test on nuScenes | 0.551 | 0.704 | 0.386 | 0.564 |
| Test on our data | 0.383 | 0.762 | 0.164 | 0.222 |
| Metrics | mAP | Car | Truck | Bus |
| Without our data | 0.417 | 0.627 | 0.292 | 0.333 |
| Concatenated with our data | 0.432 | 0.631 | 0.285 | 0.381 |
We are not aware of any previous work that fulfills the exact same task. Therefore, we compare our method to a naive inpainting baseline, which is to copy the neighboring object-free point cloud sector and paste and tile it in the masked area. The evaluation results are shown in Tab. I. Our method outperforms the baseline in all metrics, showing the effectiveness of our framework. Additionally, qualitative comparisons are shown in Fig. 6, where it is evident that our model inpaints the background based on the context, resulting in point clouds that appear natural and realistic.
V-B Object insertion
The second part of our framework is object insertion. It is designed to allow users to specify the number, type, and pose of the objects to be inserted. In Fig. 7, we show some qualitative examples of object insertion.
However, the choice of object placement will also affect the quality of the generated data and matters if we want to evaluate the generated data quantitatively. Therefore, we design different object insertion strategies when we evaluate different aspects of the object insertion algorithm. Specifically, we have two experiments. The first reveals the domain gap between the generated data and the real data, and the second showcases the value of the generated data in improving downstream tasks like object detection.
V-B1 Domain gap analysis
To rule out the impact of the choice of object placement on the realism of the generated data, we insert objects at the same pose as in the true Lidar scans. Specifically, we first remove all objects from the original dataset to obtain object-free Lidar scans. Then, we sample objects from the object library and insert them into the Lidar scan at the pose of the removed objects. In this way, we obtain a synthesized point cloud with the same pose layout of objects but with different objects. Note that while we do not change the pose layout, the pose of an inserted object is typically significantly different from the object’s original pose, making it a challenging object insertion task.
We evaluate the quality of the generated point clouds by running an object detection network VoxelNext [29] pretrained on nuScenes. The results are in Tab. II. As expected, the network performs differently on the true nuScenes dataset compared to our generated dataset. However, the results are comparable, with categories both outperforming and underperforming, suggesting a small domain gap between the two datasets.
V-B2 Value for downstream tasks
We also aim to demonstrate the practical value of our generated data in improving perception algorithms. To do this, we combine the real-world dataset with our generated dataset to fine-tune a detection network and evaluate its performance. In this experiment, we intentionally design our generated data to have different object layouts than the real data to introduce more variability into the training set. Specifically, we randomly perturb the target object poses by up to 2.5 meters and 45 degrees in yaw angle, while ensuring they remain in the ground area on the BEV map. The inserted objects are sampled from the object library. Tab. III shows that our generated data significantly improves the performance of the detection model. It shows that our synthetic data provides significant value in building more powerful algorithms.
VI Conclusions
In this paper, we contribute a novel framework for LiDAR scene editing, generating realistic synthetic data with novel object layouts in a real-world environment. We leverage generative models to fill in the missing information about both the background and the objects, and adopt a spherical voxelization model to handle Lidar projection geometry efficiently. We demonstrate the effectiveness of our framework through experiments on a large-scale self-driving dataset, nuScenes, showing that the generated LiDAR point clouds closely resemble real-world data. We show incorporating these synthetic LiDAR scans as additional training data leads to significant improvements in object detection performance in real-world data. For future work, we plan to extend the object editing task beyond vehicle-type objects. We will also use novel view synthesis to build a flexible framework, where the novel-view information may come from temporarily continuous observations when available and from prior knowledge embedded in generative models otherwise.

### v2

@@
-Generative LiDAR Editing with Controllable Novel Object Layouts
+LiDAR-EDIT: LiDAR Data Generation by Editing the Object Layouts in Real-World Scenes
 Abstract
-We propose a framework to edit real-world Lidar scans with novel object layouts while preserving a realistic background environment.
-Compared to the synthetic data generation frameworks where Lidar point clouds are generated from scratch, our framework focuses on new scenario generation in a given background environment, and our method also provides labels for the generated data.
-This approach ensures the generated data remains relevant to the specific environment, aiding both the development and the evaluation of algorithms in real-world scenarios.
-Compared with novel view synthesis, our framework allows the creation of counterfactual scenarios with significant changes in the object layout and does not rely on multi-frame optimization.
-In our framework, the object removal and insertion are supported by generative background inpainting and object point cloud completion, and the entire pipeline is built upon spherical voxelization, which realizes the correct Lidar projective geometry by construction.
-Experiments show that our framework generates realistic Lidar scans with object layout changes and benefits the development of Lidar-based self-driving systems.
+We present LiDAR-EDIT, a novel paradigm for generating synthetic LiDAR data for autonomous driving.
+Our framework edits real-world LiDAR scans by introducing new object layouts while preserving the realism of the background environment.
+Compared to end-to-end frameworks that generate LiDAR point clouds from scratch, LiDAR-EDIT offers users full control over the object layout, including the number, type, and pose of objects, while keeping most of the original real-world background.
+Our method also provides object labels for the generated data.
+Compared to novel view synthesis techniques, our framework allows for the creation of counterfactual scenarios with object layouts significantly different from the original real-world scene.
+LiDAR-EDIT uses spherical voxelization to enforce correct LiDAR projective geometry in the generated point clouds by construction.
+During object removal and insertion, generative models are employed to fill the unseen background and object parts that were occluded in the original real LiDAR scans.
+Experimental results demonstrate that our framework produces realistic LiDAR scans with practical value for downstream tasks.
+Project website with open-sourced code: https://sites.google.com/view/lidar-edit
 I Introduction
-Data is a key factor in the realization of self-driving.
-Collecting a huge amount of data with good diversity is necessary for understanding the corner cases, which are low-frequency and high-risk real-world situations.
-Both the development and the testing of self-driving systems could benefit from data with good long-tail coverages.
-However, the low-frequency nature means that it is inherently expensive to collect corner-case data, and as progress is being made on relatively easier and more frequent cases, capturing even rarer events only becomes an even more challenging task.
-That’s why the research community has a growing interest in data generation, which allows for acquiring new data at a lower cost and potentially higher diversity compared with real-world data collection.
-Early efforts of data synthesis rely on gaming engines and graphical rendering.
-This approach generates data from highly configurable virtual environments, but the scalability is limited, and the domain gap to the real data is non-negligible.
-Recent progress in generative modeling makes end-to-end data generation possible, largely improving the scalability and realism.
-However, users generally have less control over the specifics of the generated data.
-Another notable trend is novel-view synthesis from real-world sequential observations, allowing the resimulation of a real-world scenario from different viewpoints, which helps analyze and improve the algorithms.
-However, the variation flexibility of the re-simulated data is limited by the visible signal.
-In this paper, we propose to solve the task of scene editing with controllable novel object layouts, with a specific focus on Lidar scans—a widely used and reliable sensory modality in autonomous vehicles.
-This new task enables us to freely remove and insert objects with arbitrary poses in a real Lidar scan, not limited by the visible information.
-On the other hand, the generated data is associated with a real environment, building a more direct connection to real-world performance.
-An illustration of this task is shown in Fig.
+Data plays a crucial role in the advancement of self-driving technology.
+High-quality data that accurately captures the complexity and diversity of real-world driving scenes is essential for both the development and testing of autonomous systems.
+However, real-world data collection is inherently limited due to its high cost and potential safety hazards, especially when capturing long-tail corner cases, which occur infrequently and involve significant risk.
+Furthermore, real-world data simply represents a fixed recording of past events, making it inflexible for modifying specific factors to analyze their causal impact on self-driving algorithm performance.
+These challenges have motivated research focused on generating synthetic data for autonomous driving.
+Early efforts rely on gaming engines and graphical rendering.
+This approach generates data from highly configurable virtual environments, but it induces a non-negligible sim-to-real gap, and the scalability is limited.
+Recent advancements in generative models [1, 2], have enabled end-to-end data generation [3, 4], making large-scale and realistic synthetic data possible.
+However, this approach comes with a significant drawback: users have limited control over the generated data, such as precisely specifying where a newly generated car will be positioned within the scene.
+Another notable trend is novel-view synthesis from real-world sequential observations [5], allowing the rendering of a recorded scenario from different viewpoints.
+However, the variation flexibility of the rendered data is limited by the visible signal.
+In this paper, we propose a new paradigm for synthetic data generation: generative 3D scene editing with controllable novel object layouts.
+We focus on LiDAR scans—a widely used and reliable sensory modality in autonomous vehicles.
+This paradigm strikes a balance between three key factors: faithfulness, controllability, and scalability.
+Faithfulness: the generated data are only partial modifications of the real-world data, preserving the characteristics of the real environment and objects in the original data, thereby reducing the synthetic-to-real domain gap.
+Controllability: users have full control over the object layout in the generated data, including the number, type, and pose of objects, which better facilitates comparative analysis.
+Scalability: our paradigm does not rely on a physics engine, prebuilt 3D assets, or reconstruction of the whole scene.
+Instead, it is generative, facilitating efficient and scalable data creation based on vast amounts of existing real-world data.
+An illustration of our proposed method for this paradigm, LiDAR-EDIT, is shown in Fig.
 1.
-Leveraging recent progress in conditional lidar data generation and object point cloud completion, we build a lightweight framework without scene reconstruction, neural rendering, or end-to-end generation from scratch.
-We bake the prior knowledge of Lidar projective geometry into the framework through spherical voxelization so that the occlusion relationship and the scanning pattern are correct by construction.
-Experiments show that our framework enables flexible, controllable, and realistic Lidar scene editing, which is beneficial for both the development and testing of self-driving algorithms.
+Our framework first removes unwanted objects from the scene and inpaints missing regions using a generative model.
+We then insert the point clouds of objects, which are completed from their partial LiDAR scan, at user-specified poses.
+We use spherical voxelization to ensure that the occlusion and the scanning pattern follow the LiDAR projective geometry.
+Experimental results demonstrate that our framework enables flexible, controllable, and realistic LiDAR scene editing, offering significant benefits for both the development and testing of autonomous driving algorithms.
 II Related Work
-We will review the literature related to Lidar data generation.
-Most work did not directly address the same problem as targeted in this paper, but a lot of them are relevant and prepared the techniques that are useful in our problem.
+We review existing paradigms for LiDAR data generation in the literature.
 II-A LiDAR simulation
-Lidar simulation refers to the approach of building a 3D environment with 3D object assets and casting rays on the 3D environment to create point clouds through ray-surface intersection [1].
-While graphics-based simulation engines [2, 3, 4] facilitate the rendering, the requirement of pre-built 3D assets limits the scalability, and the rendering and the quality of the 3D assets both contribute to the domain gap compared with real data.
-[5] constructed the 3D assets from real-world scans to avoid manual asset creation and used learning-based postprocessing to reduce the domain gap.
-Later work [6] further investigated various Lidar phenomena that could improve the fidelity of Lidar simulators.
+LiDAR simulation creates a 3D environment with object assets and generates point clouds via ray-surface intersection.
+[6].
+Graphics-based engines [7, 8, 9] face scalability limits from pre-built 3D asset requirements, while asset quality and rendering contribute to domain gaps.
+[10] used real-world scans to generate 3D assets and applied learning-based postprocessing to reduce this gap, while [11] explored LiDAR phenomena to enhance simulator fidelity.
 II-B LiDAR novel-view synthesis
-The novel-view synthesis (NVS) of Lidar point clouds and Lidar simulation share the same conceptual procedure of 3D construction and rendering, but today’s Lidar NVS realizes these two steps in a unified learnable framework called neural rendering.
-The learned implicit geometry-appearance model and the differentiable rendering are both optimized towards reproducing more realistic synthesis, allowing a smaller domain gap and a more automatic workflow than traditional simulation [7].
-[8] modeled Lidar effects like secondary returns and ray dropping in the neural rending framework.
-[9, 10] also leveraged image sequences for additional information.
-[11, 12] explicitly the temporal dimension to account for the moving objects.
+LiDAR novel-view synthesis (NVS) and LiDAR simulation follow the same 3D construction and rendering process, but modern LiDAR NVS integrates both steps into a unified learnable framework called neural rendering.
+This approach optimizes a learned implicit geometry-appearance model through differentiable rendering for more realistic synthesis, reducing the domain gap and enabling a more automated workflow than traditional simulation [12].
+[5] incorporated LiDAR effects like secondary returns and ray dropping, while [13, 14] used image sequences for additional data.
+[15, 16] accounted for moving objects by modeling the temporal dimension.
 [17] rendered both LiDAR scans and camera images from the neural field.
-These approaches realized the playback of a real-world scenario with changed viewpoints and object states, but the pose of the observer and the objects cannot deviate too much from reality, or the rendering quality will deteriorate.
+These methods enable real-world scenario playback with altered viewpoints and object states, though significant deviations in the observer or object poses degrade rendering quality.
 II-C End-to-end LiDAR data generation
@@
 The task in this paper, LiDAR scan editing with novel object layout, is related to but different from all the tasks discussed above.
-Our output is connected to a real scene, similar to the NVS task, but we do not rely on reconstruction, which allows more flexibility in the object manipulation.
+Our output is connected to a real scene, similar to the NVS task, but we do not rely on reconstruction, which allows more flexibility in object manipulation.
 The task is generative but does not build scenes from scratch, emphasizing the controllability of object layouts.
@@
 We address the problem of generating realistic synthetic LiDAR data for autonomous driving.
-Rather than generating data completely from scratch, our approach builds upon real-world scenes, and we aim to manipulate the dynamic objects while preserving the background environment.
-The static background of a Lidar scene, denoted as , consists of stationary elements such as streets, trees, buildings, and other fixed background.
-excludes any foreground dynamic objects of interest, for example, vehicles.
-The foreground objects in the scene are represented by .
-A typical driving scene, denoted as , includes both background and foreground objects: .
-Since it is not feasible to sense the entire from a single viewpoint, we define an observation function , which transforms the scene into a partial-view Lidar point cloud: , which represents the real sensor output for an autonomous vehicle.
+Rather than generating data from scratch, we build upon real-world scenes to manipulate foreground objects while preserving the background environment.
+The static background of a LiDAR scene, , consists of stationary elements like streets, trees, and buildings, excluding foreground dynamic objects of interest such as vehicles, represented by .
+A typical driving scene includes both: .
+Since sensing the entire from a single viewpoint is infeasible, we define an observation function that transforms the scene into a partial-view point cloud: , representing the real sensor output.
 The task of LiDAR generative editing can be formally defined as follows.
@@
 IV-A Overview
-Our scene editing framework is illustrated in Fig.
+Our scene editing framework LiDAR-EDIT is illustrated in Fig.
 2.
 We first introduce the spherical voxel representation (Sec.
-IV-B), which is the foundation of our entire pipeline for the efficient modeling of the Lidar projective geometry.
-Then, we present the three components in our framework.
-The first involves removing unwanted objects from the Lidar scene (Sec.
-IV-C).
-Given an object to be removed, we identify the voxels occupied by the object’s point cloud.
-We mask these voxels and all voxels occluded by them with an inpainting mask.
-We remove points in the masked voxels and use a generative point cloud inpainting model to fill the background.
-In the second stage, we collect an object library composed of dense full-shape point clouds for every object in all Lidar scans (Sec.
-IV-D), so that we have a wide range of objects to be inserted back into the Lidar scans with flexible poses.
-We use the segmentation masks within each Lidar scene to extract partial-view point clouds of vehicles.
-These partial views are then transformed into full point clouds using a learning-based point cloud completion method.
-In the third stage, we insert new objects into the Lidar point cloud (Sec.
-IV-E).
-We start by selecting an object from the full-view object library, and specifying a desired pose for its placement.
-The dense point cloud of this object is then positioned on the ground at the specified pose.
-The resampling of the inserted objects and the occlusion are then both handled in the spherical voxel representation.
-It is worth noting that in our framework, moving an object to a different pose is treated as both removing and re-adding the same object.
-This approach helps resolve any gaps in the background and allows for significant pose changes.
+IV-B), which forms the foundation of our pipeline for efficiently modeling LiDAR projective geometry.
+Then, we present the three core components of LiDAR-EDIT.
+The first removes unwanted objects from the LiDAR scene (Sec.
+IV-C) by identifying voxels occupied by the object and voxels occluded by them.
+We mask these voxels and use a generative inpainting model to fill the background.
+In the second stage, we build an object library of dense full-shape point clouds (Sec.
+IV-D) for insertion with flexible poses.
+We extract partial-view point clouds using segmentation masks and transform them into full point clouds via learning-based completion.
+In the third stage, we insert objects into the LiDAR point cloud (Sec.
+IV-E) by selecting an object from our library and specifying a desired pose.
+The object is positioned on the ground, with resampling and occlusion handled through spherical voxelization.
+In our framework, moving an object is implemented as removing and re-adding it, which resolves background gaps and enables significant pose changes.
 IV-B Spherical Voxelization
-The spherical voxel representation is core to our entire pipeline for efficient occlusion handling and density control, as illustrated in Fig.
-3.
-Spherical voxelization discretizes the space around the origin based on three parameters : radius , azimuth , and elevation .
-This approach closely mirrors how a LiDAR scan operates, where rays are emitted from the sensor (origin) at specific angles, allowing point resampling with a pattern consistent with real LiDAR sensors.
-Voxels sharing the same azimuth and elevation values resemble a single LiDAR ray, embedding the occlusion relationship into the data structure (see Fig.
+The spherical voxel representation is core to our pipeline for efficient occlusion handling and density control (Fig.
+3).
+Spherical voxelization discretizes space around the origin using three parameters: radius , azimuth , and elevation .
+This mirrors LiDAR operation, where rays emit from the sensor at specific angles, allowing point resampling with a pattern consistent with real LiDAR sensors.
+Voxels sharing azimuth and elevation values form a single LiDAR ray, embedding occlusion relationships directly (Fig.
 3-right).
-While the range-view representation could also model the occlusion, the 2D projection loses the 3D structure.
-Preserving the 3D structure allows effective feature learning and convenient point cloud manipulation in our pipeline.
+While range-view representation could model occlusion, it loses 3D structure in 2D projection.
+Preserving this 3D structure enables effective feature learning and point cloud manipulation.
 IV-C Object removal
-The object removal has two steps: identifying the area affected by the removal of an object and generating background points in this area that are consistent with the surrounding environment.
-We call these two steps de-occlusion masking and background inpainting, respectively.
+The object removal stage involves two steps: identifying the area affected by object removal and generating consistent background points in this area.
+We term these steps de-occlusion masking and background inpainting.
 IV-C1 De-occlusion masking
-Given the spherical voxelization of the input, a de-occlusion mask refers to a binary mask over the voxels, where the voxels occupied or occluded by the object to be removed are labeled positive.
-The occupancy of the masked voxels needs to be predicted because of the scene change (object removed) and the visibility change (occlusion removed).
-We first mask the voxels occupied by the object point cloud.
-Then, all voxels sharing the same azimuth and elevation indices with larger radius indices than the occupied voxels are also masked (see Fig.
+Given spherical voxelization and the object to be removed, the de-occlusion mask highlights voxels whose occupancy needs prediction after removal—specifically, voxels occupied or occluded by the object.
+We first mask voxels occupied by the object, then mask all voxels sharing the same azimuth and elevation with larger radius indices than the occupied voxels (see Fig.
 3-right).
 IV-C2 Background inpainting
-The background inpainting can be viewed as a conditional generation problem.
-We recast the UltraLidar framework [16] for this task.
-UltraLidar can be viewed as a Lidar version of MaskGIT [17], which is a generative model for images with transformer-based multi-step autoregression in a discrete latent space.
+We approach background inpainting as a conditional generation problem using a method inspired by the UltraLidar framework [19], which is a LiDAR adaptation of MaskGIT[20] that employs transformer-based multi-step autoregression in discrete latent space.
 The training has two stages, as depicted in Fig.
 4.
-First, we train a VQ-VAE [23] to encode raw point clouds into a latent feature map with discrete latent codes.
-We use a bird’s-eye-view (BEV) 2D latent map to represent the encoded point cloud.
-Notice that our BEV uses the azimuth-radius coordinate from the spherical voxelization, meaning that the voxels along the elevation axis are compressed to a single point in the latent feature map.
-It is slightly different from the physical top-down projection and also different from the Cartesian voxelization used in UltraLidar.
-We use the original point cloud scans to learn this encoder.
-The second step is to learn a generative model in the latent space for background inpainting.
-Similar to MaskGIT [17], a bi-directional transformer is trained to map an incomplete BEV latent feature map to a complete latent feature map, which can be decoded to a point cloud.
-The feature prediction is cast as a classification problem because the feature space is discrete.
-The prediction is iterative.
-We use a mask to keep track of which tokens in the BEV feature map need to be predicted.
-In every iteration, the transformer predicts the masked tokens from the known ones, but only a subset of the predicted tokens with the top classification confidence is preserved, and they will be treated as known tokens in the next iteration.
-The other tokens remain masked and need to be predicted in the following iterations until the feature map is complete.
-The BEV mask is initialized from the voxel mask that is used to remove points from the input point cloud.
-In training, we apply the voxel mask on random object-free voxels, so that the transformer is trained to predict only the background for the masked tokens.
-At inference, we use the de-occlusion voxel mask introduced in Sec.
+First, we train a VQ-VAE [26] to encode raw point clouds into a discrete latent feature map.
+We represent this using a bird’s-eye-view (BEV) 2D map with azimuth-radius coordinates, compressing elevation-angle-axis voxels to single latent points.
+This differs from physical top-down projection and UltraLidar’s Cartesian voxelization.
+We use the original point cloud scans to train this encoder.
+The second step learns a generative model in latent space for background inpainting.
+Following MaskGIT [20], we train a bi-directional transformer to map incomplete BEV feature maps to complete ones, which decode to point clouds.
+Feature prediction is treated as a classification problem of the discrete latent codes with iterative processing.
+A mask tracks which tokens need prediction, with each iteration preserving only sampled high-confidence predictions as known tokens for subsequent iterations.
+The BEV mask initializes from the voxel mask used to remove points.
+During training, we apply the mask to random object-free voxels, training the transformer to predict only background for masked tokens.
+At inference, we use the de-occlusion mask from Sec.
 IV-C1.
-While we follow the UltraLidar framework [16], the object removal in UltraLidar is realized by replacing the token at the object position with a background discrete feature, which needs to be manually specified and does not explicitly inpaint the areas occluded by the removed objects.
-Our object-removal module resolves both problems.
+LiDAR-EDIT inherits UltraLidar’s two-stage framework [19], while tailoring the system to generate realistic backgrounds after object removal by locating occluded areas via spherical voxelization and using background-only training.
+In comparison, UltraLidar simply replaces object tokens with manually-specified discrete background features without guaranteeing complete background inpainting.
 IV-D Object library construction
@@
 We choose the point cloud as the shape representation of objects for two reasons.
-First, we can resample points from the shape conveniently in the spherical voxel representation and do not need expensive ray-casting on meshes or volumetric rendering on implicit fields.
+First, we can resample points from the shape conveniently through spherical voxelization without ray-casting on meshes or volumetric rendering on implicit fields.
 Second, point cloud completion techniques are well-established to facilitate full-shape generation.
-We applied a pretrained model of the state-of-the-art point cloud completion network, AnchorFormer [24], in our framework.
+We apply a pretrained model of the state-of-the-art point cloud completion network, AnchorFormer [27], in our framework.
 IV-E Object insertion
@@
 We assume that a user only specifies the x-y coordinate and the yaw angle of the desired insertion pose, and the insertion algorithm needs to figure out the rest of the degrees of freedom based on the existing background point cloud.
-We further assume zero pitch and roll angles as are commonly assumed in self-driving datasets [25, 26].
+We further assume zero pitch and roll angles, as commonly assumed in self-driving datasets [28, 29].
 Therefore, we are left with height to be determined.
@@
 Now, the inserted object has a pattern and density that are consistent with the overall LiDAR scan.
-In the end, we resolve the occlusion.
-For each coordinate with at least one occupied voxel, an occlusion mask is applied to all voxels with radius , where is the smallest radius of the occupied voxels in the 1D ray.
+In the end, we resolve all occlusions using spherical projection: for each coordinate with at least one occupied voxel, an occlusion mask is applied to all voxels with radius , where is the smallest radius of the occupied voxels in the 1D ray.
 All the masked voxels are set to empty.
 V Experiments and results
-This paper aims at a novel Lidar editing task that is different from the literature.
-Therefore, it is not straightforward to apply a well-established evaluation protocol and compare with existing work.
-We will separate the experiments into two major components, i.e., object removal and object insertion, so that we can validate the effectiveness of each step and show that the overall framework provides a practical solution for Lidar scene editing.
-We conduct experiments on the nuScenes-LidarSeg [25] dataset, which is composed of 40,000 point clouds sampled from 1,000 driving scenes.
-The point clouds are collected by a 32-beam Velodyne HDL32E Lidar.
-We use the official training and validation splits.
-All results are trained on the Lidar scans in the training set and evaluated using scans in the validation set.
-We focus on the manipulation of vehicle-type objects, including cars, trucks, and buses.
-The specific use of the dataset in each experiment is explained in the regarding sections.
+This paper introduces a novel LiDAR editing approach distinct from all existing methods.
+Therefore, it is not straightforward to use a well-established evaluation protocol or compare LiDAR-EDIT with previous works.
+We divide the experiments into two major components, object removal and object insertion, showing our framework generates realistic LiDAR scenes.
+We conduct experiments on the nuScenes-LidarSeg [28] dataset, consisting of 40,000 point clouds from 1,000 driving scenes, captured by a 32-beam Velodyne HDL32E LiDAR.
+Using the official training and validation splits, all models are trained on the training set of size 28,130, and evaluated on the validation set of size 6,019.
+The dataset’s specific use in each experiment is detailed in the relevant sections.
+In all experiments, we use spherical voxelization of resolution within the range of along the radius, azimuth and elevation axes respectively.
 V-A Object removal and background inpainting
-The goal of this experiment is to test that our approach can inpaint the background points when objects are removed.
-The tricky thing is that we do not have the ground truth background behind an object.
-Therefore, we create artificial occlusion masks over the background area in a Lidar scan and ask our model to recover the masked area.
-In this way, we have the ground truth background to facilitate evaluation.
-In order to make the artificial masks similar to the ones caused by the occlusion of real objects, we create the masks from the bounding boxes of objects in nuScenes.
-This process is illustrated in Fig.
-5.
-We calculate the average size of the 3D bounding boxes of all vehicles in nuScenes.
-Then, we put the bounding box at a fixed distance, with the heading orthogonal to the viewing direction, so that the box occludes the most azimuth range.
-We rotate the bounding box around the ego vehicle till the bounding box falls into an object-free sector of the Lidar scan to be masked.
-We pick the distance as 10 meters to create a substantial occlusion.
-We evaluate the performance of our background inpainting using statistical metrics and perceptual metrics.
-For the statistical metrics, we leverage the Maximum-Mean Discrepancy (MMD) and the Jensen-Shannon Divergence (JSD).
-Similar to [16], we construct a 2D histogram along the ground plane and use the voxel occupancy instead of the number of points for the bin count.
-However, our evaluation has two differences from that in [16].
-First, our 2D histogram is in the azimuth-radius coordinate.
-Second, we manually set the count for all bins that are not in the generated area to zero and re-normalize the histogram to sum to one, because we only generate the point cloud for a masked area instead of generating the whole point cloud from scratch.
-For the perceptual metrics, we use the Frechet Sparse Volume Distance (FSVD) and the Frechet Point-based Volume Distance (FPVD).
+We design two experiments to evaluate the quality of our inpainted background and to verify that our pipeline can generate object-free point clouds after object removal and background inpainting, respectively.
+The first experiment evaluates LiDAR-EDIT’s ability to inpaint realistic background after vehicle removal.
+Since ground truth background behind real objects is not available, we create artificial de-occlusion masks over background areas and task our model with reconstructing these masked points, giving us ground-truth for evaluation.
+To create masks resembling real object occlusions, we use bounding boxes from nuScenes objects (Fig.
+5).
+We calculate the average size of all ground truth vehicle bounding boxes, then place a box of the average size at a fixed 10-meter distance with heading orthogonal to viewing direction to maximize azimuth occlusion.
+We rotate this box until it falls within an object-free sector of the LiDAR scan.
+We evaluate our generated background against ground truth using both statistical and perceptual metrics.
+Statistical metrics quantifythe global structural differences between the two sets of background poins, while perceptual metrics assess differences in the latent feature space, capturing both global and local structures.
+For statistical metrics, we use Maximum-Mean Discrepancy (MMD) and Jensen-Shannon Divergence (JSD).
+Following [19], we construct a 2D histogram along the ground plane, using voxel occupancy instead of point count for bin values.
+Our evaluation differs in two ways: (1) our 2D histogram is in azimuth-radius coordinates, and (2) we zero-out bins outside the generated area and re-normalize the histogram, as we generate only a masked region rather than an entire point cloud.
+For perceptual metrics, we use Frechet Sparse Volume Distance (FSVD) and Frechet Point-based Volume Distance (FPVD).
 Following [18], FSVD is evaluated using MinkowskiNet [30], and FPVD is evaluated using SPVCNN [31], both pretrained on nuScenes.
@@
 | Ours | 0.169 | 0.173 | 0.584 | 6.61 |
-| Metrics | mAP | Car | Truck | Bus |
-| Test on nuScenes | 0.551 | 0.704 | 0.386 | 0.564 |
-| Test on our data | 0.383 | 0.762 | 0.164 | 0.222 |
-| Metrics | mAP | Car | Truck | Bus |
-| Without our data | 0.417 | 0.627 | 0.292 | 0.333 |
-| Concatenated with our data | 0.432 | 0.631 | 0.285 | 0.381 |
-We are not aware of any previous work that fulfills the exact same task.
-Therefore, we compare our method to a naive inpainting baseline, which is to copy the neighboring object-free point cloud sector and paste and tile it in the masked area.
-The evaluation results are shown in Tab.
-I.
-Our method outperforms the baseline in all metrics, showing the effectiveness of our framework.
-Additionally, qualitative comparisons are shown in Fig.
-6, where it is evident that our model inpaints the background based on the context, resulting in point clouds that appear natural and realistic.
+| Data generation method | Ours | UltraLidar [19] |
+| Number of objects detected per frame | 2.6 | 15.5 |
+As no prior work addresses this exact task, we compare our method to a naive inpainting baseline, which copies and tiles a neighboring object-free point cloud sector into the masked area.
+Using the nuScenes validation set for evaluation, Tab.
+I shows that our method outperforms the baseline across all metrics, demonstrating its effectiveness.
+Fig.
+6 provides qualitative comparisons, highlighting how our model inpaints the background based on the context, producing more natural and realistic point clouds.
+In the second experiment, we use de-occlusion masks from ground-truth object bounding boxes to generate object-free point clouds.
+We then use an object detection model, VoxelNext [32], to count objects in the point clouds, and the results are in Tab.
+II.
+Ideally, the count should be zero; however, a small nonzero value is observed due to false positives, which is expected.
+For reference, UltraLidar’s unconditionally generated point clouds, which are expected to contain objects, yield an average of 15.5 detected objects.
+| Training data | Testing data | mAP | car | non-car |
+| nuScenes | Ours | 0.340 | 0.785 | 0.290 |
+| nuScenes | nuScenes | 0.361 | 0.672 | 0.326 |
+| Ours + nuScenes | nuScenes | 0.368 | 0.672 | 0.334 |
 V-B Object insertion
 The second part of our framework is object insertion.
-It is designed to allow users to specify the number, type, and pose of the objects to be inserted.
+It is designed to allow users to freely specify the number, type, and pose of the objects to be inserted.
+In this experiment, we only manipulate the car category because AnchorFormer’s [27] point cloud completion is less reliable on other categories, but our framework is agnostic of the choice of completion models.
 In Fig.
 7, we show some qualitative examples of object insertion.
-However, the choice of object placement will also affect the quality of the generated data and matters if we want to evaluate the generated data quantitatively.
+The choice of object placement will also affect the quality of the generated data and matters if we want to evaluate the generated data quantitatively.
 Therefore, we design different object insertion strategies when we evaluate different aspects of the object insertion algorithm.
 Specifically, we have two experiments.
-The first reveals the domain gap between the generated data and the real data, and the second showcases the value of the generated data in improving downstream tasks like object detection.
+The first reveals the domain gap between the generated synthetic data and the real data, and the second showcases the value of our synthetic data in improving autonomous driving downstream tasks such as object detection.
 V-B1 Domain gap analysis
 To rule out the impact of the choice of object placement on the realism of the generated data, we insert objects at the same pose as in the true LiDAR scans.
-Specifically, we first remove all objects from the original dataset to obtain object-free Lidar scans.
-Then, we sample objects from the object library and insert them into the Lidar scan at the pose of the removed objects.
-In this way, we obtain a synthesized point cloud with the same pose layout of objects but with different objects.
-Note that while we do not change the pose layout, the pose of an inserted object is typically significantly different from the object’s original pose, making it a challenging object insertion task.
-We evaluate the quality of the generated point clouds by running an object detection network VoxelNext [29] pretrained on nuScenes.
-The results are in Tab.
-II.
-As expected, the network performs differently on the true nuScenes dataset compared to our generated dataset.
-However, the results are comparable, with categories both outperforming and underperforming, suggesting a small domain gap between the two datasets.
-V-B2 Value for downstream tasks
-We also aim to demonstrate the practical value of our generated data in improving perception algorithms.
-To do this, we combine the real-world dataset with our generated dataset to fine-tune a detection network and evaluate its performance.
-In this experiment, we intentionally design our generated data to have different object layouts than the real data to introduce more variability into the training set.
+Specifically, we first remove all objects and inpaint the background.
+Then, we sample random objects from the object library and insert them into the LiDAR scan at the pose of the removed objects.
+In this way, we obtain a synthesized point cloud that maintains the original pose layout of objects while replacing the objects with different ones.
+We apply this procedure to the entire nuScenes validation set, producing a synthetic validation set of the same size as the original.
+We assess the quality of our generated synthetic point clouds using the object detector VoxelNext [32], pretrained on nuScenes.
+We evaluate VoxelNext on both the (1) real and (2) synthetic nuScenes validation sets.
+In the first two rows of Tab.
+III, we can see that the mAP (mean average precision over all categories) remains similar when tested on the original validation set compared to our synthetic data, indicating a small domain gap.
+The car category’s higher AP shows that our inserted object point clouds have a small domain gap with the original ones, but they may also be too simple due to the lack of ray-drop modeling, indicating a direction for future improvements.
+The slightly lower APs of other categories are mostly due to the increased false positives caused by the background inpainting, which is helpful as we will see in Sec.
+V-B2.
+V-B2 Values on autonomous driving downstream tasks
+We aim to demonstrate the practical value of our generated data in improving perception algorithms.
+We compare two VoxelNext models, trained from scratch, using two different training procedures.
+The first model is trained only on real nuScenes data for 40 epochs.
+The second model is first pretrained on our synthetic training dataset (details provided below) for 15 epochs; then trained on real nuScenes data for an additional 40 epochs, following the same training schedule as the first model.
+In this experiment, we design our synthetic data to have different object layouts than the real data to introduce more variability.
 Specifically, we randomly perturb the target object poses by up to 2.5 meters and 45 degrees in yaw angle, while ensuring they remain in the ground area on the BEV map.
 The inserted objects are sampled from the object library.
+We apply this procedure to the entire nuScenes training set, generating a synthetic training set of the same size for pretraining.
 Tab.
-III shows that our generated data significantly improves the performance of the detection model.
-It shows that our synthetic data provides significant value in building more powerful algorithms.
+III shows that pretraining on our generated data (third row) yields superior performance compared to training exclusively on real nuScenes (second row).
+This shows that our synthetic data provides significant value in building more powerful algorithms.
+Interestingly, the improvement mostly comes from non-car categories by allowing the model to better suppress false positives.
+The inserted car objects, which are slightly easier to detect than the real ones, did not change the performance of the final model, but we expect it to improve with more detailed LiDAR modeling like ray-drop.
 VI Conclusions
-In this paper, we contribute a novel framework for LiDAR scene editing, generating realistic synthetic data with novel object layouts in a real-world environment.
-We leverage generative models to fill in the missing information about both the background and the objects, and adopt a spherical voxelization model to handle Lidar projection geometry efficiently.
-We demonstrate the effectiveness of our framework through experiments on a large-scale self-driving dataset, nuScenes, showing that the generated LiDAR point clouds closely resemble real-world data.
-We show incorporating these synthetic LiDAR scans as additional training data leads to significant improvements in object detection performance in real-world data.
-For future work, we plan to extend the object editing task beyond vehicle-type objects.
-We will also use novel view synthesis to build a flexible framework, where the novel-view information may come from temporarily continuous observations when available and from prior knowledge embedded in generative models otherwise.
+This paper introduces a novel paradigm for LiDAR data generation by editing the object layouts in real-world LiDAR scans.
+Following this paradigm, our LiDAR-EDIT framework generates data with a small domain gap to the real world and gives users high controllability.
+Experiments show practical value of the generated data in downstream autonomous driving tasks.
+Incorporating ray-drop effects, optimizing object shape models, and automating the pose layout generation are interesting improving directions for future work.